# What’s in an embedding? Would a rose by any embedding smell as sweet?

### v2

@@ Abstract
Large Language Models (LLMs) are often criticized for lacking true ”understanding” and an ability to ”reason” with their knowledge, being seen merely as advanced autocomplete systems. We believe that this perspective might be missing an important insight. We suggest that LLMs do develop a kind of empirical ”understanding” that is ”geometry”-like, which seems quite sufficient for a range of applications in NLP, computer vision, coding assistance, etc. However, this ”geometric” understanding, built from incomplete and noisy data, makes them unreliable, difficult to generalize, and lacking in inference capabilities and explanations, similar to the challenges faced by heuristics-based expert systems decades ago.
To overcome these limitations, we suggest that LLMs should be integrated with an ”algebraic” representation of knowledge that includes symbolic AI elements used in expert systems. This integration aims to create large knowledge models (LKMs) that not only possess ”deep” knowledge grounded in first principles, but also have the ability to reason and explain, mimicking human expert capabilities. To harness the full potential of generative AI safely and effectively, a paradigm shift from LLMs to the more comprehensive LKMs is needed.
@@ I Introduction
“What’s in a name? That we call a rose by any other name would smell as sweet.” These immortal words from Shakespeare’s Romeo and Juliet resonate in the context of large language models. One might ask: What’s in an embedding? Would a rose by any embedding still smell as sweet? In other words, does an embedding capture all the properties of an entity comprehensively?
Related to this is the important question: Do LLMs truly ”understand” the meaning of the words and sentences they generate as humans do, or are they merely performing autocomplete without any ”understanding?”
@@
So, when one asks Georgette for an explanation about how she reached her conclusions, her answer might be something like: ”Points A and B are on the circle because they are! And I can see that they are on opposite ends and that the angle between them is $80^{\circ}$. There is no need to explain any of this!” She cannot explain her conclusions to anyone who does not have access to her representation.
@@
Therefore, even in highly precise disciplines such as mathematics, we can have representations that are incomplete. This dichotomy also suggests the exciting possibility that there could exist other representations in high-dimensional spaces that are ”richer” in capturing and revealing important mathematical relationships. This possibility reminds us of the almost magical abilities of Srinivasa Ramanujan, the Indian mathematical genius, for whom scrolls containing the most complicated mathematics would unfold in his dreams [2]. Ramanujan seems to have had access to an abundantly rich representation of mathematics in his mind that was more ”geometric” or ”visual” in nature, in the spirit of our discussion above. Therefore, many complex and startling results were obvious to him because he ”saw” them to be true. For him, the new mathematical results did not require proof because they are like points A and B in Fig. 2. They obviously lie on the circumference and, therefore, needed no proof or explanation.
@@ III Knowledge representation in LLMs
Studying the evolution of GPT-1 to GPT-3.5 in the context of our discussion above, we find that this is similar to the representation of a circle with more and more noisy data points, as in Fig. 4. Figs. 1-2 represent the perfect circle, the one with infinite noise-free points. On the other hand, Fig. 4 represents different imperfect, incomplete, and noisy approximations of a circle. Generally speaking, GPT-1 is like Fig. 4a, whereas GPT-3.5 is like Fig. 4c. Just as interpolations or extrapolations made using Fig. 4a could lead to incorrect results, the sparse data in GPT-1 made its performance relatively weak. Again, as in Fig. 4c, the much denser data in GPT-3.5 resulted in a much better performance.
@@
In the real world, the knowledge equivalent of the ”perfect circle” is not available for most practical situations. So, what we do is develop useful approximations by training LLMs on enormous amounts of data that tend to be inherently incomplete and uncertain. This results in LLMs developing an internal ”geometric” representation of their noisy training data similar to Fig. 4.
@@
Recent articles from researchers in Anthropic AI and Open AI shed some light on the internal representations of commercial-grade LLMs [3, 4, 5]. The Anthropic AI team examined one of their LLMs, Claude 3 Sonnet, using a method called ”dictionary learning” to discover patterns in the activation of neuron combinations when Claude was asked to discuss specific topics. Approximately 10 million of these patterns or features were identified.
@@
These results resonate with our hypothesis and analysis that the LLMs are using a ”geometry”-like internal representation in a high-dimensional embedding space rather than an ”algebraic” one. In the circle example, the geometric representation readily revealed which points are spatially close to each other, which are further apart, what the subtended angles are, etc., to make reasoning easier. Similarly, in the high-dimensional ”meanings” space represented by the vector embeddings, Claude’s internal representation captured the meanings of words and phrases and their relative distances in a ”geometry”-like manner, making it easier to do sophisticated ”reasoning,” such as analogies and metaphors. Claude ”reasons” without actually doing the ”reasoning” but simply ”looking up.” So, Claude ”understands” its knowledge in the sense that Georgette understands a circle.
The important question here is the following: What is the ”algebraic” representation of Claude’s knowledge? This is possible for simple geometric objects like circles, but is it even possible for complex large-scale data? If not, can we at least develop useful ”analytical” approximations?
@@
In a biological visual system, features are extracted, assembled, and interpreted at hierarchical levels of detail. Similarly, we believe that the LLMs similarly perform ”picture”-like processing and interpretations of the tokens. In this sense, the LLMs seem to possess a ”mind’s eye”, as it were. Many queries that require considerable effort of searching and reasoning for us would seem obvious and easy, requiring no reasoning or search for them. Like Georgette, the LLMs do not ”reason” because they don’t have to! They do not ”explain” because there is nothing to ”explain”!
@@ IV From LLMs to LKMs for Science and Engineering
Although the connectionist camp seems to have won the latest battle for ideas, we believe that, in the long run, AI will be determined by the successful integration of both the symbolic and connectionist ideas. We need the symbolic paradigm to capture first-principles-based mechanistic ”deep” knowledge and reasoning, and connectionist representations for data-driven empirical knowledge [6, 7, 8, 9, 10]. .
@@ V More is Different
In 1972, the physics Nobel laureate Philip Anderson published a highly influential paper entitled ”More is Different”[15]. In it, he observes:
@@
Likewise, the ultra-large language models, such as GPT-3.5 and beyond, are not mere autocomplete engines, but they have new emergent capabilities that require creating a new conceptual framework similar to the transition from Newtonian mechanics to statistical mechanics. As noted, in statistical mechanics, $F=ma$ is happening in molecular collisions, but the overall physics is much more than and much different from what is implied by Newton’s laws. Similarly, in GPT-3.5 and beyond, autocomplete is happening at the individual token level, but the overall behavior is much more and much different. More is different, indeed!
In conclusion, generative AI has opened up unimagined possibilities in all aspects of human endeavor. It is poised to have a great impact on many aspects of science and engineering. However, to harness its potential safely and effectively, we need to go beyond large language models to large knowledge models, which deeply incorporate fundamental knowledge and human expertise. This requires integrating the ”geometric” and ”algebraic” representations of the world in a new mathematical framework of large knowledge models. Obviously, nature has figured out how to do this using deep neural networks in our brains. Our challenge is to do the same in generative AI systems.

### v3

@@ Abstract
Large Language Models (LLMs) are often criticized for lacking true ”understanding” and the ability to ”reason” with their knowledge, being seen merely as autocomplete systems. We believe that this assessment might be missing a nuanced insight. We suggest that LLMs do develop a kind of empirical ”understanding” that is ”geometry”-like, which seems adequate for a range of applications in NLP, computer vision, coding assistance, etc. However, this ”geometric” understanding, built from incomplete and noisy data, makes them unreliable, difficult to generalize, and lacking in inference capabilities and explanations, similar to the challenges faced by heuristics-based expert systems decades ago.
To overcome these limitations, we suggest that LLMs should be integrated with an ”algebraic” representation of knowledge that includes symbolic AI elements used in expert systems. This integration aims to create large knowledge models (LKMs) that not only possess ”deep” knowledge grounded in first principles, but also have the ability to reason and explain, mimicking human expert capabilities. To harness the full potential of generative AI safely and effectively, a paradigm shift is needed from LLM to more comprehensive LKM.
@@ I Introduction
’What’s in a name? That we call a rose by any other name would smell as sweet.” These immortal words from Shakespeare’s Romeo and Juliet resonate in the context of large language models. One might ask: What’s in an embedding? Would a rose by any embedding still smell as sweet? In other words, does an embedding capture all the properties of an entity comprehensively?
Related to this is the important question: Do LLMs truly ”understand” the meaning of the words and sentences they generate as humans do, or are they ”stochastic parrots” [1] merely performing autocomplete without any ”understanding?”
@@
So, when one asks Georgette for an explanation about how she reached her conclusions, her answer might be something like: ”Points A and B are on the circle because they are! And I can see that they are on opposite ends. There is no need to explain any of this!” She cannot explain her conclusions to anyone who does not have access to her representation.
@@
Therefore, even in highly precise disciplines such as mathematics, we can have representations that are incomplete. This dichotomy also suggests the exciting possibility that there could exist other representations in high-dimensional spaces that are ”richer” in capturing and revealing important mathematical relationships. This possibility reminds us of the almost magical abilities of Srinivasa Ramanujan, the Indian mathematical genius, for whom scrolls containing the most complicated mathematics would unfold in his dreams [3]. Ramanujan seems to have had access to an abundantly rich representation of mathematics in his mind that was more ”geometric” or ”visual” in nature, in the spirit of our discussion above. Therefore, many complex and startling results were obvious to him because he ”saw” them to be true.
@@ III Knowledge representation in LLMs
The evolution of GPT-1 to GPT-3.5 is similar to the representation of a circle with more and more noisy data points, as in Fig. 4. Figs. 1-2 represent the perfect circle, the one with infinite noise-free points. On the other hand, Fig. 4 represents different imperfect, incomplete, and noisy approximations of a circle. Therefore, for the aforementioned questions about points A & B and C & D one can get erroneous and stochastic answers depending on the approximation. Generally speaking, GPT-1 is like Fig. 4a, whereas GPT-3.5 is like Fig. 4c. As in Fig. 4c, the much denser data in GPT-3.5 resulted in much better performance.
@@
In the real world, the equivalent of the ”perfect circle” is not available in most practical situations. So, what we do is develop useful approximations by training LLMs on enormous amounts of data that tend to be inherently incomplete and uncertain. This results in LLMs developing a noisy ”geometric” approximation of the world.
@@
Recent articles from researchers in Anthropic AI [4, 5] and Open AI [6] shed some light on the internal representations of commercial grade LLMs. The Anthropic AI team examined one of their LLMs, Claude 3 Sonnet, using a method called ”dictionary learning” to discover patterns in the activation of neuron combinations when Claude was asked to discuss specific topics. Approximately 10 million of these patterns or features were identified.
@@
These results resonate with our hypothesis and analysis that the LLMs are using a ”geometry”-like internal representation in a high-dimensional embedding space rather than an ”algebraic” one. In the circle example, the geometric representation readily revealed which points are spatially close to each other, which are further apart, what the subtended angles are, etc., to make reasoning easier. Similarly, in the high-dimensional ”meanings” space represented by the vector embeddings, Claude’s internal representation captured the meanings of words and phrases and their relative distances in a ”geometry”-like manner, making it easier to do sophisticated ”reasoning,” such as analogies and metaphors. Claude ”reasons” without actually doing the ”reasoning” but simply ”looking up.” We suggest that Claude ”understands” its knowledge in the sense that Georgette understands a circle.
An important question here is the following: What is the ”algebraic” representation of Claude’s knowledge? This is possible for simple geometric objects like circles, but is it even possible for complex large-scale data? If not, can we at least develop useful ”analytical” approximations?
@@
In a biological visual system, features are extracted, assembled, and interpreted at hierarchical levels of detail. Similarly, we believe that the LLMs similarly perform ”picture”-like processing and interpretations of the tokens. In this sense, the LLMs seem to possess a ”mind’s eye”, as it were. An LLM is more like Cyclops than a stochastic parrot. Many queries that require considerable effort of searching and reasoning for us would seem obvious and easy, requiring no reasoning or search for them. Like Georgette, the LLMs do not ”reason” because they don’t have to! They do not ”explain” because there is nothing to ”explain”!
@@ IV From LLMs to LKMs for Science and Engineering
Although the connectionist camp seems to have won the latest battle for ideas, we believe that, in the long run, AI will be determined by the successful integration of both the symbolic and connectionist ideas. We need the symbolic paradigm to capture first-principles-based mechanistic ”deep” knowledge and reasoning, and connectionist representations for data-driven empirical knowledge [7, 8, 9, 10, 11, 15].
@@ V More is Different
In 1972, the physics Nobel laureate Philip Anderson published an influential paper entitled ”More is Different”[17]. In it, he observes:
@@
Likewise, the ultra-large language models, such as GPT-3.5 and beyond, are not mere autocomplete engines, but they have new emergent capabilities that require creating a new conceptual framework similar to the transition from Newtonian mechanics to statistical mechanics. As noted, in statistical mechanics, $F=ma$ is happening in molecular collisions, but the overall physics is much more than and much different from what is implied by Newton’s laws. Similarly, in GPT-3.5 and beyond, autocomplete is happening at the individual token level, but the overall behavior is much more and much different. They may not have developed a human-like understanding of their domain, but they seem to have acquired a different kind of understanding and intelligence. It may be incomplete and flawed, but we do not yet know enough about their internals (or, for that matter, our own internals) to categorically deny their ”understanding” and ”intelligence”.
In conclusion, generative AI has opened up unimagined possibilities in all aspects of human endeavor. However, to harness its potential safely and effectively, we need to go beyond large language models to large knowledge models, which deeply incorporate fundamental knowledge and human expertise. This requires integrating the ”geometric” and ”algebraic” representations of the world in a new mathematical framework of large knowledge models. Obviously, nature has figured out how to do this using deep neural networks in our brains. Our challenge is to do the same in generative AI systems.